# Bayesian Estimation and Tuning-Free Rank Detection for Probability Mass Function Tensors

### v1

@@ VI Conclusion
In this work, we have presented a Bayesian PMF estimation approach that simultaneously allows the rank (model order) of the joint PMF tensor to be inferred from observed data. Our approach only requires a single training run in order to infer the rank, obviating the need for computationally intensive cross-validation procedures typical of classical model order selection techniques. Moreover, we have developed a scalable version of the PMF estimation approach which leverages stochastic optimization techniques, enabling efficient handling of massive datasets with minimal performance loss. Numerical simulations on synthetic data have demonstrated the validity of our approaches in terms of computational efficiency, robustness to missing data, tuning-free rank inference, and scalability. Additionally, results from a real-world experiment demonstrate that our proposed algorithms outperform state-of-the-art techniques in both accuracy and computational efficiency, further highlighting their practical applicability.
@@
$\mathbb{E}[\log p(\bm{Y}\,|\bm{\Theta})]$ $=$ (56)
$\mathbb{E}\Big{[}\log p\big{(}\bm{Y}\,\big{|}\,\bm{Z},\{\bm{A}_{n}\}_{n=1}^{N}\big{)}\Big{]}+\mathbb{E}\Big{[}\log p(\bm{Z}\,|\,\bm{\lambda})\Big{]}$
$+\mathbb{E}\Big{[}\log p(\bm{\lambda})\Big{]}+\sum_{n=1}^{N}\sum_{r=1}^{R}\mathbb{E}\Big{[}\log p(\bm{a}_{n,r})\Big{]},$
@@
$\mathbb{E}\Big{[}\log p\big{(}\bm{Y}\,\big{|}\,\bm{Z},\{\bm{A}_{n}\}_{n=1}^{N}\big{)}\Big{]}+\mathbb{E}\Big{[}\log p(\bm{Z}\,|\,\bm{\lambda})\Big{]}$
$=\sum_{t=1}^{T}\sum_{r=1}^{R}\mathbb{E}[z_{r,t}]\Big{(}\mathbb{E}\big{[}\log\lambda_{r}\big{]}+\sum_{n=1}^{N}\mathbb{E}\big{[}\log a_{n,r,y_{n,t}}\big{]}\Big{)}$ (57)
@@
$+(\alpha_{n,r}-1)\sum_{i_{n}=1}^{I_{n}}\mathbb{E}[\log a_{n,r,i_{n}}]$ (59)
Solving the expectations using (32) and (37) yields the first three terms of (39). Similarly, the second term on the right-hand side of (38) is given by (cf. (17))
@@
$+(\widetilde{\alpha}_{n,r}-1)\sum_{i_{n}=1}^{I_{n}}\mathbb{E}[\log a_{n,r,i_{n}}]$ (63)
after which the expectations are solved via (32) and (37) to yield the last three terms of (39).

### v2

@@ VI Conclusion
We presented a novel Bayesian framework for joint parameter and model order (rank) estimation of low-rank probability mass function (PMF) tensors. Unlike existing tensor-based PMF estimation approaches, which require the rank to be heuristically pre-selected or determined through computationally expensive validation or information criteria, the proposed method infers the rank directly from the observed data during inference. This is accomplished through a Bayesian model that exploits Dirichlet priors to enforce probability simplex constraints and to promote sparsity in the latent components, thereby enabling irrelevant rank-one terms to be pruned without ad hoc thresholding or repeated training.
The resulting VB-PMF algorithm yields closed-form posterior updates with deterministic convergence and offers computational efficiency while eliminating the need for cross-validation. Experiments using synthetic data demonstrate that the proposed method converges to the true tensor rank as the number of samples increases, and achieves a higher estimation accuracy and a better robustness to missing data than likelihood-based and coupled tensor factorization approaches. Furthermore, experiments on real classification and recommendation datasets demonstrate its practical effectiveness. On these tasks, VB-PMF equals or exceeds the performance of classical models like random forests and biased matrix factorization while offering an interpretable and versatile model based on the low-rank components of the estimated joint PMF.
@@
${\color[rgb]{0,0,0}\mathbb{E}[\log p(\bm{Y},\bm{\Theta})]}$ $=$ (43)
$\hskip-50.00008pt\mathbb{E}\Big[\log p\big(\bm{Y}\,\big|\,\bm{Z},\{\bm{A}_{n}\}_{n=1}^{N}\big)\Big]+\mathbb{E}\Big[\log p(\bm{Z}\,|\,\bm{\lambda})\Big]$
$\hskip-40.00006pt+\mathbb{E}\Big[\log p(\bm{\lambda})\Big]+\sum_{n=1}^{N}\sum_{r=1}^{R}\mathbb{E}\Big[\log p(\bm{a}_{n,r})\Big],$
@@
$\hskip-60.00009pt\mathbb{E}\Big[\log p\big(\bm{Y}\,\big|\,\bm{Z},\{\bm{A}_{n}\}_{n=1}^{N}\big)\Big]+\mathbb{E}\Big[\log p(\bm{Z}\,|\,\bm{\lambda})\Big]$
$\hskip-60.00009pt=\sum_{t=1}^{T}\sum_{r=1}^{R}\mathbb{E}[z_{r,t}]\Big(\mathbb{E}\big[\log\lambda_{r}\big]+\sum_{n=1}^{N}\mathbb{E}\big[\log a_{n,r,y_{n,t}}\big]\Big)$ (44)
@@
$+(\alpha_{n,r}-1)\sum_{{\color[rgb]{0,0,0}i}=1}^{I_{n}}\mathbb{E}[\log a_{n,r,{\color[rgb]{0,0,0}i}}]$ (46)
Calculating the expectations via (32) and (37) yields the first three terms of (39). Similarly, the second term on the right-hand side of (38) is given by (cf. (17))
@@
$+(\widetilde{\alpha}_{n,r}-1)\sum_{{\color[rgb]{0,0,0}i}=1}^{I_{n}}\mathbb{E}[\log a_{n,r,{\color[rgb]{0,0,0}i}}]$ (50)
after which the expectations are computed via (32) and (37) to yield the last three terms of (39).